OpenStudy (anonymous):

2 numbers are first added together and then multiplied together. The numbers added together = 18 and the numbers multiplied together = 45. What are the 2 numbers? The numbers are 15 and 3

OpenStudy (anonymous):

is this like a proof by chance?

OpenStudy (anonymous):

if its not then i guess its just 15+ 3 = 18 and 15*3 = 45

OpenStudy (id21):

You need to solve a equation system. \[x+y=18\] and \[x \times y =45\]

OpenStudy (id21):

\[x=18-y\] and put it in second equation: \[x \times y = 45; (18-y)y=45\] \[y^2-18y+45=0\] \[y _{1,2}=15;3\] put it in first equation and find the x \[x _{1,2}=18-y _{1,2}=3;15\] Answer: x=3 when y=15 or x=15 when y=3